MONT 103N – Analyzing Environmental Data Paper 1 – Connections Reaction Paper Due Date: Friday, February 24

General Information

The first paper assignment for the semester will involve tying together:

• Chapter 8 from Sherry Turkle’s Alone Together

• Your reactions to Sherry Turkle’s talk on Thursday, February 16

• Your own reactions to and opinions about the Connections event on February 17 - 18.

Your essay should be 3-5 double-spaced pages in length (but don’t worry if you go a bit over or are a bit short – the length is not the main point). Prepare your essay using Microsoft Word or equivalent software, and submit your work to me by email to:

                           jlittle@holycross.edu

This is due no later than 5:00pm on Friday, February 24.

The General Topic

Based on your reading of the chapter from Alone Together, Sherry Turkle’s talk, our class discussions, and discussions with your classmates during the Connections event or at any other time, what do you think about the concerns expressed by Turkle that everpresent communications technology is affecting the quality of our interactions with others? Starting with what you presumably know best, what is your “life-mix” like? Do you feel the sense of being “always on” that Turkle describes – the sense of being overwhelmed by texts, calls, emails and other communications that you must deal with? Or are the main issues and challenges you face something different? Is choosing to communicate almost exclusively in those ways motivated by a desire to be “never alone but always in control” (Turkle, p. 157) ? Do you feel that way yourself? Do you know anyone who spends a major part of his or her “life mix” in role-playing games online (like “Rolo” and “Jade” – Turkle, p. 159)? What do you think about that choice? Of course, being like the “young fish,” it can be difficult to realize that the “water” is there if you don’t know anything else. Did you take this opportunity to explore different ways of interacting with the world seriously, or did you “blow it off?” (Be honest!) Has this event helped you to understand that it is possible to choose whether or not to use this technology and also to choose the extent to which we use it? If so, what aspects were most meaningful for you? Do you think you will change as a result? If not, did you see any value in this exercise?

More Details

This will not be a “research paper” and it will not be necessary to consult any sources other than the chapter from Alone Together and your own personal notes from February 16 - 18. But the “they say/I say” framework will still be extremely useful to keep in mind (especially if you want to say that Turkle’s concerns about what communications technology is doing to our interpersonal relationships are overstated, or irrelevant, or wrong in some other way, but also if you end up largely agreeing with her point of view). A good paper should definitely have a specific point of view about the questions above, and the first thing you will want to do is to identify what it is you want to say and how you will tie everything together.

As always, write in a reasonably formal style incorporating correct grammar, spelling, and usage. Your essay should be well-structured with paragraphs corresponding to the major points in your outline, arranged in a logical sequence. (Feel free to use some of the “templates” from “They Say/I Say” to help structure things – no footnotes are necessary for that.) But your job here is mainly to figure out exactly what you think about these questions and to present those thoughts in a coherent and well-structured essay.

If you want to use quotation(s) from the Turkle chapter, use the following style:

People’s personalities develop at different rates and as Turkle says, ”[n]o handle cranks, no gear turns to move us from one stage of life to another” (Turkle, p. 152).

It is permissible but not necessary to use other sources. Direct quotations or paraphrases of other information should follow the following guidelines:

1) Identify direct quotations with quotation marks

2) All direct quotations and paraphrases of arguments found in other sources should be identified by notes in parentheses giving information where the quote can be found within your source,

3) List all your sources in a References section at the end of your paper:

a) For books, give the author(s), the title, city where the publisher is located, the publisher, and year of publication.

b) For magazine or newspaper articles, give the author(s), the title of the article, the title of the magazine or newspaper, the date of publication, and the starting and final page number of the specific article.

c) For web sites, give an author (if you can determine that), the full URL (web address) and the date you accessed it (since information on the web often changes!)